On February 1, 2020, Pat Weaver Inc. (PWI) issued 7%, $1,600,000 bonds for $1,900,000. PWI retired all of these bonds on January 1, 2021, at 104. Unamortized bond premium on that date was $166,400. How much gain or loss should be recognized on this bond retirement?

Multiple Choice

  • $112,000 gain.

  • $0 gain.

  • $102,400 gain.

  • $133,000 gain.
Answer #1

Carrying value of bonds = Face value+Unamortized premium = 1600000+166400 = 1766400

Retirement value = 1600000*1.04 = 1664000

Gain = 1766400-1664000 = 102400

So answer is c) $102400 Gain